OpenStudy (anonymous):

How many solutions does the equation have? A. one solution B. infinite solutions C. no solution

OpenStudy (anonymous):

@coolman500

OpenStudy (anonymous):

I would help you but you think it's edgy to aesthetisize and revision a dark history of white supremacist violence

OpenStudy (anonymous):

What you mean by that

OpenStudy (anonymous):

@welshfella

OpenStudy (anonymous):

@AJ01

OpenStudy (anonymous):

I'm here sorry about that

OpenStudy (anonymous):

what do you think

OpenStudy (anonymous):

it ok so i tried it on my calculator and whenever i press equals it answers the question and dont get to type the whole question

OpenStudy (anonymous):

okay well what do you think for the question not the answer

OpenStudy (anonymous):

a+5=1/5(5a+25) 5(a+5)=5a+25 5a+25=5a+25

OpenStudy (anonymous):

okay I know that but what do you think the answer is to the question

OpenStudy (anonymous):

oh a

OpenStudy (anonymous):

no not A

OpenStudy (anonymous):

@AJ01 Don't give direct answers

OpenStudy (anonymous):

http://openstudy.com/code-of-conduct

OpenStudy (anonymous):

so what do you think it's not A we know that

OpenStudy (anonymous):

c

OpenStudy (anonymous):

there is always a solution so what does that mean

OpenStudy (anonymous):

b

OpenStudy (anonymous):

yup good job

OpenStudy (anonymous):

remember there is always a solution so never would it be C